The following monthly data are available for the Winfast Company and its only product:

Unit selling price $36

Unit variable expenses $28

Total fixed expenses $51,000

Actual sales of the month of March  7,000

The margin of safety for the company during March was:
Answer #1

Break even unit = Fixed cost/Contribution margin per unit = 51000/(36-28) = 6375 Units

Margin of safety = Actual Sales-Break even sales = 7000-6375 = 625 Units